Science Operations

Category: Announcements

PEPSI PFU SX Vignetting – Resolved

Sep 15, 2026

9/14/26: During summer shutdown 2026 restart, it was discovered that the PEPSI SX PFU side was suffering from severe vignetting despite attempts to realign with the M3 on-sky.  The cause was tracked down to index-matching oil in between some of …

Read Now

Restart 2026

Aug 31, 2026

Overview of SSD26 Telescope Activities
The SX and DX primary and tertiary mirrors were washed during SSD26.
The observing computers were upgraded to AlmaLinux 9. This upgrade introduces new procedures for remote desktop access. Users should refer to the Remote …

Read Now

TMS Installation Delayed – Resolved

Aug 28, 2026

8/28/26: Due to last minute staffing commitments, the TMS system will not be available until approximately late September.  However the LBCs can still be used although more frequent re-collimations will be necessary.
9/15/26: TMS system has been reinstalled, science release …

Read Now

PEPSI Chamber Pressure Control Issue – Resolved

Aug 27, 2026

8/9/26: The PEPSI chamber pressure dropped down and began following atmospheric pressure.  The issue was traced to a proportional valve that was always open and exhausting air to the outlets at the top of the chamber.  Replacement parts have been …

Read Now

MODS Controller Upgrade

Apr 09, 2026

MODS  Controller Upgrade – MODS1 released for shared-risk observations
update: Apr 28th
MODS1 is now released for shared-risk observations. The release of MODS2 is being held back by an issue with the slitmask cassette which is being investigated.
Unlike with …

Read Now

LUCI1 Detector Focus Failure – Resolved

Mar 27, 2026

LUCI1 Detector Focus Repaired, Returned for Operations (May 4, 2026)
LUCI1 has been released for SL/ESM science. Verification sets showed that the passive flexure tables from Sept 2025 looked good, with the spots remaining within ~2 pixels over the full …

Read Now

LBT Users Meeting

Feb 24, 2026

There will be an LBT Users Meeting on Tuesday, March 3 at 8am Arizona time over Zoom. Please contact the sciops team for login information.

Read Now

TMS PC Down – Resolved

Feb 17, 2026

Feb. 12, 2026: The TMS system recently experienced a hardware issue with the GPU in the PC controlling the system.  A replacement GPU is currently being sourced, but for the time TMS will be offline.
March 27, 2026: We have …

Read Now

MODS2 Mask Cassette Issues – Resolved

Jun 28, 2025

On the night of June 26 UT when moving the 2.4″ longslit into the focal plane, the mask failed to insert into the focal plane despite the software thinking so.  When the instrument was opened, the grabber was found to …

Read Now

MOS Mask Milling Machine Issue – Resolved

Apr 10, 2025

April 11, 2025: Problem with the oxygen regulator was found and resolved.  The MOS mask milling machine has been returned to operation.
April 9, 2025: The laser for the MOS mask milling machine recently stopped working.  Currently all mask milling …

Read Now